OpenStudy (anonymous):

How do you solve the equation sin(x) = 1/2 for all values of x in between 0 and 2pi, including 0 but not 2pi?

OpenStudy (anonymous):

Pick out all the angles with a sine of 1/2 from the unit circle.

OpenStudy (zzr0ck3r):

memorize the unit circle and the "common" angles

OpenStudy (turingtest):

\[\sin\theta=x\iff\theta=\sin^{-1}x\]so to solve for x, or whatever you call the angle, you must take the inverse sine of both sides there is no simple process for this, so you either have to be familiar with the unit circle, manipulate the expression with identities, or use a calculator depending on the situation

OpenStudy (zzr0ck3r):

I geuarnte your teacher wants you to know them

OpenStudy (anonymous):

So basically, pi/6 and 5pi/6 according to the Unit Circle?